OpenStudy (anonymous):

suppose that f(x)=-x-7 and g(x)= x-12 solve for f(x)=0 solve for g(x)=0 solve for f(x)=g(x)

OpenStudy (anonymous):

i know f(x)0=-7 g(x)0=-12 i just dont know the third one

OpenStudy (anonymous):

f(x)=-x-7 g(x)= x-12 solve for f(x)=g(x) -x-7=x-12 rearrange and solve. good luck

OpenStudy (anonymous):

not sure still

OpenStudy (anonymous):

-x-7=x-12 add x to both sides -7 = 2x-12 add 12 to both sides 5 = 2x